OpenStudy (anonymous):

if the angle of elevation of the sun is 75 degrees 30' how tall is post that casts shadow 1.5 ft long

OpenStudy (anonymous):



OpenStudy (anonymous):

sin (75.5) = x/1.5 (x in feet) x = 1.5*sin (75.5)